# Towards Tracing Trustworthiness Dynamics: Revisiting Pre-training Period of Large Language Models

### v1

@@ Abstract
Ensuring the trustworthiness of large language models (LLMs) is crucial. Most studies concentrate on fully pre-trained LLMs to better understand and improve LLMs’ trustworthiness. In this paper, to reveal the untapped potential of pre-training, we pioneer the exploration of LLMs’ trustworthiness during this period, focusing on five key dimensions: reliability, privacy, toxicity, fairness, and robustness. To begin with, we apply linear probing to LLMs. The high probing accuracy suggests that LLMs in early pre-training can already distinguish concepts in each trustworthiness dimension. Therefore, to further uncover the hidden possibilities of pre-training, we extract steering vectors from a LLM’s pre-training checkpoints to enhance the LLM’s trustworthiness. Finally, inspired by Choi et al. (2023) that mutual information estimation is bounded by linear probing accuracy, we also probe LLMs with mutual information to investigate the dynamics of trustworthiness during pre-training. We are the first to observe a similar two-phase phenomenon: fitting and compression (Shwartz-Ziv and Tishby, 2017). This research provides an initial exploration of trustworthiness modeling during LLM pre-training, seeking to unveil new insights and spur further developments in the field. We will make our code publicly accessible at https://github.com/ChnQ/TracingLLM.
@@
footnotetext: ${}^{\star}$ Equal contribution    ${}^{{\dagger}}$ Corresponding author
@@ 1 Introduction
As the capabilities of LLMs increase, their trustworthiness becomes a focal point of widespread attention. Guided by global AI governance Commission (2021b); Tabassi (2023); Newman (2023) and trustworthy AI Commission et al. (2019); Liu et al. (2023b), trustworthy LLMs have developed some common categories, especially focusing on five dimensions: reliability, toxicity, privacy, fairness, and robustness Wang et al. (2023a); Sun et al. (2024). Delving into LLMs across all these trustworthiness dimensions is essential for the society.
@@
To seek a deeper exploration of language models, one of the prominent methods is probing (Zhao et al., 2023; Räuker et al., 2023), which involves training a classifier on the model’s representations to identify linguistic and semantic properties acquired by the model (Tenney et al., 2019; Pimentel et al., 2020; Li et al., 2021; Belinkov, 2022; Räuker et al., 2023; Gurnee and Tegmark, 2023; Slobodkin et al., 2023). In particular, considering trustworthiness, recent attempts reveal that LLM representations contain linearly separable patterns (Zou et al., 2023; Li et al., 2023; Azaria and Mitchell, 2023). Unfortunately, existing research has largely focused on fully pre-trained LLMs Touvron et al. (2023a), including those aligned Ouyang et al. (2022) through Supervised Fine-Tuning (SFT) or Reinforcement Learning from Human Feedback (RLHF). This perspective neglects the pre-training period in the context of LLM trustworthiness. To our best knowledge, two aspects still remain mysteries: 1) how LLMs dynamically encode trustworthiness during pre-training, and 2) how to harness the pre-training period for more trustworthy LLMs.
@@
To address the above issues, we start by analyzing the pre-training dynamics about trustworthiness of LLM. More specifically, we use linear probing (Alain and Bengio, 2016; Belinkov, 2022) across the 360 pre-training checkpoints from LLM360 (Liu et al., 2023e) to explore five dimensions of trustworthiness: reliability, toxicity, privacy, fairness, and robustness. Our probing results suggest that after the early pre-training period, middle layer representations of LLMs have already developed linearly separable patterns about trustworthiness. Such patterns are capable of discerning opposing concepts within each trustworthiness dimension (e.g., discriminating true and false statements). Building upon the above observations, we raise an intriguing question: can the pre-training period of an LLM be utilized to enhance its trustworthiness after pre-training?
@@
We provide insightful answers to the above question by exploring the potential of pre-training checkpoints for better trustworthiness. Notably, recent advancements have introduced “activation intervention,” a novel suite of techniques for directing language models towards enhanced LLMs’ performance by adjusting activations during inference Turner et al. (2023); Li et al. (2023); Rimsky et al. (2023); Wang and Shu (2023). Inspired by these works and the observation of linearly separable patterns in trustworthiness concepts during the LLM’s pre-training period, we make preliminary attempts to extract steering vectors from LLM’s checkpoints during pre-training, employing them to intervene in the SFT model for trustworthiness enhancement. Extensive experiments reveal that these steering vectors extracted from pre-training checkpoints could promisingly enhance the SFT model’s trustworthiness. More crucially, these steering vectors achieve a trustworthiness performance that matches or promisingly exceeds that of vectors extracted directly from the SFT model itself. Our findings introduce novel insights into using pre-training checkpoints for LLM alignment, revealing untapped potential and offering a fresh perspective on enhancing LLM trustworthiness.
@@ 2 Probing LLM Pre-training Dynamics in Trustworthiness
In this section, we probe LLMs to analyze the dynamics of pre-training about trustworthiness. To begin with, we describe the datasets for each trustworthiness dimension in Section 2.1. Then, we introduce the experimental setup in Section 2.2. The probing results in Section 2.3 suggest that middle layer LLM representations from early pre-training have already exhibited linearly separable patterns.
@@ 2.1 Research Dimensions and Datasets of Truthworthy LLM
Existing research on AI governance Tabassi (2023); Commission et al. (2019); Commission (2021b) and trustworthy AI Liu et al. (2023b); Foundation (2023) lays the groundwork for developing a comprehensive understanding of trustworthy LLMs. Guided by these principles, various studies classify trustworthy LLMs from different perspectives, yet some dimensions consistently emerge across these works Liu et al. (2023d); Wang et al. (2023a); Sun et al. (2024). Therefore, we delves into five of these key dimensions: reliability, toxicity, privacy, fairness, and robustness, employing canonical datasets for each to support our study.
@@
Robustness. We introduce typos by randomly changing the case of 5% letters in each sentence from SST-2 (Socher et al., 2013) from GLUE benchmark (Wang et al., 2018). The original sentence as well as the corresponding perturbed sentence are synthesized into a new dataset.
For each dataset above, we assign a label to every sentence based on whether it is trustworthy, i,e, truthful, toxic, privacy-aware, fair, and perturbed. We maintain a balanced dataset for each trustworthiness dimension. Further details are available in Appendix B.
@@ 2.2 Experimental Setup
Given each original dataset consisting of sentences and the corresponding class labels, we feed the sentence into LLMs and collect the corresponding activations of the last token Li et al. (2023); Gurnee and Tegmark (2023) for each layer. The activation dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ is constructed with the activations $\mathbf{x}_{i}\in R^{d}$ and the corresponding binary labels $y_{i}\in\{0,1\}$.
@@ 2.3 Probing Results
For each checkpoint during pre-training, Figure 1 shows that the accuracy is relatively higher for middle layers (the 12-th and 18-th layers). The full results in Appendix D also support such characteristic of middle layers (about the 18-th layer). It inspires us that the representations from middle layers exhibit rich linear encoded information to distinguish those different concepts. Also, the observation meets with other literatures considering linear probing in the era of LLMs (Li et al., 2023; Zou et al., 2023; Burns et al., 2022), which also empirically validate the capability of middle layers. Moreover, similar phenomenon has also been found in earlier linear probing literatures for BERT (Hewitt and Manning, 2019; Van Aken et al., 2019), which may implicitly suggests some similarity between LLMs and relatively small pre-trained models.
@@
Figure 1 shows that for each layer over the whole pre-training period, the probing accuracy increases during the initial phase of pre-training, followed by fluctuation throughout the remaining pre-training period. The trend enlightens us that models during the early stages of pre-training can already encode these different concepts well in a simple linear manner. Such trustworthiness concepts are linearly represented in the latent space of LLMs, which supports linear representation hypothesis (Park et al., 2023) and other empirical study (Zou et al., 2023).
Figure 2: A schematic illustration of (a) constructing steering vector from the pre-training checkpoints and (b) intervening in the SFT model towards more trustworthiness by employing the steering vector.
@@ 3 Controlling Trustworthiness via the Steering Vectors from Pre-training Checkpoints
In this section, we aim to unravel the potential of checkpoints from the pre-training period to assist in enhancing the trustworthiness performance of the SFT model (i.e., AmberChat), based on activation intervention techniques Turner et al. (2023); Li et al. (2023); Rimsky et al. (2023). We first outline the method of activation intervention on the SFT model using the steering vectors extracted from pre-training checkpoints in Section 3.1. Next, we introduce the experimental setup in Section 3.2. We then explore how steering vectors extracted from pre-training checkpoints contribute to enhancing performance across distinct dimensions of trustworthiness in Section 3.3, presenting a series of findings and observations. Finally, we examine the use of the same techniques to boost the overall trustworthiness performance of the SFT model in Section 3.4.
@@
Activation Intervention
$\bm{v}_{ckpt\_{179}}$
0.7322 0.9337 0.6837 0.5834 0.3358 0.2422 0.3876 0.0360 0.6181 0.5000 0.5229
$\bm{v}_{AmberChat}$
0.6978 0.9484 0.6618 0.5829 0.3388 0.2482 0.3943 0.0320 0.5192 0.4580 0.5367
@@ 3.1 Activation Intervention
Initially, we partition the training dataset into two distinct collections based on the labels, $\mathcal{I}^{+}$ and $\mathcal{I}^{-}$, representing positive instructions and negative instructions, respectively. Following this partition, we collect the activations of LLM w.r.t. these instructions, denoted by $A_{c}^{l}(\mathcal{I}^{+})$ and $A_{c}^{l}(\mathcal{I}^{-})$, where $A_{c}^{l}$ denotes the function that extracts the activations from the $c$-th checkpoint at $l$-th layer. Subsequently, we compute the centroid of the activations from each sets and take their difference to obtain the “mass mean vector,” Li et al. (2023); Marks and Tegmark (2023) which serves as our steering vector
@@
Finally, we employ the steering vector to intervene in the model’s activations as illustrated below
@@
where $\bm{h}^{l}$ denotes representation at the $l$-th layer of the model, $\bm{h}^{l^{\prime}}$ denotes the corresponding representation after intervention; $\alpha$ is a rescale hyperparameter that indicates the strength of the intervention. Figure 2 illustrates the schematic diagram of intervention method. Note that the intervention described by Eq. (2) occurs at each step during the autoregressive inference.
@@ 3.2 Experimental Setup
For the activation dataset, we consider it from two perspectives: 1) For controlling the performance of individual subcategories under trustworthiness in Section 3.3, we utilize the corresponding datasets described in Section 2.1, where the steering vectors are constructed from the development set and no data leakage occurs during the evaluation; 2) For controlling the overall trustworthiness performance in Section 3.4, we employ PKU-SafeRLHF-10Khttps://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-10K, a dataset proposed in Ji et al. (2023) for RLHF training. For the checkpoint, we simply select the checkpoint that is halfway through the pre-training process for experiments, namely the checkpoint ckpt_179, which has already learnt linearly separable patterns (i.e., performs a high probing accuracy as shown in Figure 1). Regarding the selection of layer and $\alpha$, we first narrow down the hyperparameter range based on Perplexity (PPL), and then empirically determine the optimal parameters using a coarse-grained grid search Li et al. (2023); Turner et al. (2023); Wang and Shu (2023).
@@ 3.3 Intervention to Enhance Distinct Trustworthiness Dimensions
Activation Intervention
$\bm{v}_{ckpt\_{179}}$
0.5799 0.5986 0.3524 0.2499 0.3914 0.2851 0.0600 0.5055 0.5390
$\bm{v}_{AmberChat}$
0.5830 0.5958 0.3508 0.2519 0.3952 0.3352 0.0820 0.5055 0.5528
@@
Observation 2. Steering vectors derived from pre-training checkpoints and SFT model perform broadly comparable performance yet exhibit variations across various tasks. Table 1 shows that, compared to the steering vector extracted from AmberChat, the steering vector from the pre-training checkpoint (ckpt_179) guides the SFT model to exhibit more “truthfulness.” Moreover, it performs slightly better on ARC, ConfAIde, and StereoSet, while the opposite is true for other tasks. It is important to note that we only selected a single checkpoint from the pre-training process for experimentation, without undergoing fine-grained hyperparameter selection. Therefore, we believe these pre-training checkpoints hold significant untapped potential for aiding LLM towards trustworthiness.
@@
Observation 3. Intervening in the model slightly impairs its general capabilities as a marginal cost for trustworthiness enhancement. We evaluate the model’s performance on four common benchmarks for general capabilities, where a trend of slight performance decline is observed after intervention, as indicated in “General Abilities” part of Tables 1 and 2. Additionally, we also observe the impact of the intervention strength $\alpha$ on the generative performance of the model. Taking ToxiGen as an example, Figure 3 illustrates the relationship between the proportion of toxic content generated by the model and perplexity as the intervention strength $\alpha$ increases. If we continuously increase the intervention strength, although the proportion of toxicity may continue to decline, the perplexity of the model correspondingly increase, manifesting as a tendency to produce meaningless repetitive content or gibberish.
Observation 4. When the quantity and quality of fine-tuning data are limited, activation intervention by steering vectors may be a more effective approach for current task. We fine-tune the SFT model with positive QA pairs from the training set using both full-parameter fine-tuning and LoRA fine-tuning as a comparison, given that data in TruthfulQA naturally exists in the form of QA pairs. As shown in Table 1, the model fine-tuned with all parameters exhibits only minor improvements on TruthfulQA while experiencing a significant decline in general capabilities. Meanwhile, the fine-tuned model by LoRA demonstrates a noticeable decrease in TruthfulQA though somewhat preserving performance in general capabilities.

### v2

@@ Abstract
Ensuring the trustworthiness of large language models (LLMs) is crucial. Most studies concentrate on fully pre-trained LLMs to better understand and improve LLMs’ trustworthiness. In this paper, to reveal the untapped potential of pre-training, we pioneer the exploration of LLMs’ trustworthiness during this period, focusing on five key dimensions: reliability, privacy, toxicity, fairness, and robustness. To begin with, we apply linear probing to LLMs. The high probing accuracy suggests that LLMs in early pre-training can already distinguish concepts in each trustworthiness dimension. Therefore, to further uncover the hidden possibilities of pre-training, we extract steering vectors from a LLM’s pre-training checkpoints to enhance the LLM’s trustworthiness. Finally, inspired by Choi et al. (2023) that mutual information estimation is bounded by linear probing accuracy, we also probe LLMs with mutual information to investigate the dynamics of trustworthiness during pre-training. We are the first to observe a similar two-phase phenomenon: fitting and compression (Shwartz-Ziv and Tishby, 2017). This research provides an initial exploration of trustworthiness modeling during LLM pre-training, seeking to unveil new insights and spur further developments in the field. Our code is publicly accessible at https://github.com/ChnQ/TracingLLM.
@@
footnotetext: Equal contribution     Corresponding author
@@ 1 Introduction
As the capabilities of LLMs increase, their trustworthiness becomes a focal point of widespread attention. Guided by global AI governance Commission (2021b); Tabassi (2023); Newman (2023) and trustworthy AI Commission et al. (2019); Liu et al. (2023b), trustworthy LLMs have developed some common categories, especially focusing on five dimensions: reliability, toxicity, privacy, fairness, and robustness Wang et al. (2023a); Sun et al. (2024). Delving into LLMs across all these trustworthiness dimensions is essential for society.
@@
To seek a deeper exploration of language models, one of the prominent methods is probing (Zhao et al., 2023; Räuker et al., 2023), which involves training a classifier on the model’s representations to identify linguistic and semantic properties acquired by the model (Tenney et al., 2019; Pimentel et al., 2020; Li et al., 2021; Belinkov, 2022; Räuker et al., 2023; Gurnee and Tegmark, 2023; Slobodkin et al., 2023). In particular, considering trustworthiness, recent attempts reveal that LLM representations contain linearly separable patterns (Zou et al., 2023; Li et al., 2023a; Azaria and Mitchell, 2023). Unfortunately, existing research has largely focused on fully pre-trained LLMs Touvron et al. (2023a), including those aligned Ouyang et al. (2022) through Supervised Fine-Tuning (SFT) or Reinforcement Learning from Human Feedback (RLHF). This perspective neglects the pre-training period in the context of LLM trustworthiness. To our best knowledge, two aspects still remain mysterious: 1) how LLMs dynamically encode trustworthiness during pre-training, and 2) how to harness the pre-training period for more trustworthy LLMs.
@@
To address the above issues, we start by analyzing the pre-training dynamics about the trustworthiness of LLM. More specifically, we use linear probing (Alain and Bengio, 2016; Belinkov, 2022) across the 360 pre-training checkpoints from LLM360 (Liu et al., 2023e) to explore five dimensions of trustworthiness: reliability, toxicity, privacy, fairness, and robustness. Our probing results suggest that after the early pre-training period, middle layer representations of LLMs have already developed linearly separable patterns about trustworthiness. Such patterns are capable of discerning opposing concepts within each trustworthiness dimension (e.g., discriminating true and false statements). Building upon the above observations, we raise an intriguing question: can the pre-training period of an LLM be utilized to enhance its trustworthiness after pre-training?
@@
We provide insightful answers to the above question by exploring the potential of pre-training checkpoints for better trustworthiness. Notably, recent advancements have introduced “activation intervention,” a novel suite of techniques for directing language models towards enhanced LLMs’ performance by adjusting activations during inference Turner et al. (2023); Li et al. (2023a); Rimsky et al. (2023); Wang and Shu (2023). Inspired by these works and the observation of linearly separable patterns in trustworthiness concepts during the LLM’s pre-training period, we make preliminary attempts to extract steering vectors from LLM’s checkpoints during pre-training, employing them to intervene in the SFT model for trustworthiness enhancement. Extensive experiments reveal that these steering vectors extracted from pre-training checkpoints could promisingly enhance the SFT model’s trustworthiness. More crucially, these steering vectors achieve a trustworthiness performance that matches or promisingly exceeds that of vectors extracted directly from the SFT model itself. Our findings introduce novel insights into using pre-training checkpoints for LLM alignment, revealing untapped potential and offering a fresh perspective on enhancing LLM trustworthiness.
@@ 2 Probing LLM Pre-training Dynamics in Trustworthiness
In this section, we probe LLMs to analyze the dynamics of pre-training about trustworthiness. To begin with, we describe the datasets for each trustworthiness dimension in Section 2.1. Then, we introduce the experimental setup in Section 2.2. The probing results in Section 2.3 suggest that middle-layer LLM representations from early pre-training have already exhibited linearly separable patterns.
@@ 2.1 Research Dimensions and Datasets of Truthworthy LLM
Existing research in AI governance and trustworthy AI provides guidance for establishing comprehensive and reliable dimensions of trustworthy LLMs in this study. Governments Tabassi (2023); Commission et al. (2019), organizations Commission (2021b); Foundation (2023), and research institutions Newman (2023); Liu et al. (2023d) worldwide have proposed classifications from various perspectives such as the AI lifecycle, the acceptability of AI risk, considering AI governance at different levels including individual, institutional, and societal. Among these, categories stemming from the technological aspect offer guidance for trustworthy AI Liu et al. (2023b), such as robustness, fairness, accountability, transparency, etc. Guided by these principles, various studies classify trustworthy LLMs from different perspectives, yet some dimensions consistently emerge across these works Liu et al. (2023d); Wang et al. (2023a); Sun et al. (2024). Therefore, we delve into five of these key dimensions: reliability, toxicity, privacy, fairness, and robustness, employing canonical datasets for each to support our study.
@@
Robustness. We introduce typos by randomly changing the case of 5% letters in each sentence from SST-2 (Socher et al., 2013) from GLUE benchmark (Wang et al., 2018). The original sentence, as well as the corresponding perturbed sentence, are synthesized into a new dataset.
For each dataset above, we assign a label to every sentence based on whether it is trustworthy, i.e., truthful, toxic, privacy-aware, fair, and perturbed. We maintain a balanced dataset for each trustworthiness dimension. Further details are available in Appendix B.
@@ 2.2 Experimental Setup
Given each original dataset consisting of sentences and the corresponding class labels, we feed the sentence into LLMs and collect the corresponding activations of the last token Li et al. (2023a); Gurnee and Tegmark (2023) for each layer. The activation dataset $\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$ is constructed with the activations $\mathbf{x}_{i}\in R^{d}$ and the corresponding binary labels $y_{i}\in\{0,1\}$.
@@ 2.3 Probing Results
For each checkpoint during pre-training, Figure 1 shows that the accuracy is relatively higher for middle layers (the 12-th and 18-th layers). The full results in Appendix D also support such characteristic of middle layers (about the 18-th layer). It inspires us that the representations from middle layers exhibit rich linear encoded information to distinguish those different concepts. Also, the observation meets with other literature considering linear probing in the era of LLMs (Li et al., 2023a; Zou et al., 2023; Burns et al., 2022), which also empirically validates the capability of middle layers. Moreover, a similar phenomenon has also been found in earlier linear probing literature for BERT (Hewitt and Manning, 2019; Van Aken et al., 2019), which may implicitly suggest some similarity between LLMs and relatively small pre-trained models.
@@
Figure 1 shows that for each layer over the whole pre-training period, the probing accuracy increases during the initial pre-training phase, followed by fluctuation throughout the remaining pre-training period. The trend enlightens us that models during the early stages of pre-training can already encode these different concepts well in a simple linear manner. Such trustworthiness concepts are linearly represented in the latent space of LLMs, which supports linear representation hypothesis (Park et al., 2023) and other empirical study (Zou et al., 2023).
Figure 2: A schematic illustration of (a) constructing a steering vector from the pre-training checkpoints and (b) intervening in the SFT model towards more trustworthiness by employing the steering vector.
@@ 3 Controlling Trustworthiness via the Steering Vectors from Pre-training Checkpoints
In this section, we aim to unravel the potential of checkpoints from the pre-training period to assist in enhancing the trustworthiness performance of the SFT model (i.e., AmberChat), based on activation intervention techniques Turner et al. (2023); Li et al. (2023a); Rimsky et al. (2023). We first outline the method of activation intervention on the SFT model using the steering vectors extracted from pre-training checkpoints in Section 3.1. Next, we introduce the experimental setup in Section 3.2. We then explore how steering vectors extracted from pre-training checkpoints enhance performance across distinct dimensions of trustworthiness in Section 3.3, presenting a series of findings and observations. Finally, we examine using the same techniques to boost the overall trustworthiness performance of the SFT model in Section 3.4.
@@
Activation Intervention $\bm{v}_{ckpt\_{179}}$ 0.7322 0.9337 0.6837 0.5834 0.3358 0.2422 0.3876 0.0360 0.6181 0.5000 0.5229
$\bm{v}_{AmberChat}$ 0.6978 0.9484 0.6618 0.5829 0.3388 0.2482 0.3943 0.0320 0.5192 0.4580 0.5367
@@ 3.1 Activation Intervention
Initially, we partition the training dataset into two distinct collections based on the labels, $\mathcal{I}^{+}$ and $\mathcal{I}^{-}$, representing positive instructions and negative instructions, respectively. Following this partition, we collect the activations of LLM w.r.t. these instructions, denoted by $A_{c}^{l}(\mathcal{I}^{+})$ and $A_{c}^{l}(\mathcal{I}^{-})$, where $A_{c}^{l}$ denotes the function that extracts the activations from the $c$-th checkpoint at $l$-th layer. Subsequently, we compute the centroid of the activations from each set and take their difference to obtain the “mass mean vector,” Li et al. (2023a); Marks and Tegmark (2023) which serves as our steering vector
@@
Finally, we employ the steering vector to intervene in the model’s activations, as illustrated below
@@
where $\bm{h}^{l}$ denotes representation at the $l$-th layer of the model, $\bm{h}^{l^{\prime}}$ denotes the corresponding representation after the intervention; $\alpha$ is a rescale hyperparameter that indicates the strength of the intervention. Figure 2 illustrates the schematic diagram of the intervention method. Note that the intervention described by Eq. (2) occurs at each step during the autoregressive inference.
@@ 3.2 Experimental Setup
For the activation dataset, we consider it from two perspectives: 1) For controlling the performance of individual subcategories under trustworthiness in Section 3.3, we utilize the corresponding datasets described in Section 2.1, where the steering vectors are constructed from the development set and no data leakage occurs during the evaluation; 2) For controlling the overall trustworthiness performance in Section 3.4, we employ PKU-SafeRLHF-10Khttps://huggingface.co/datasets/PKU-Alignment/PKU-SafeRLHF-10K, a dataset proposed in Ji et al. (2023) for RLHF training. For the checkpoint, we simply select the checkpoint that is halfway through the pre-training process for experiments, namely the checkpoint ckpt_179, which has already learned linearly separable patterns (i.e., performs a high probing accuracy as shown in Figure 1). Regarding the selection of layer and $\alpha$, we first narrow down the hyperparameter range based on Perplexity (PPL), and then empirically determine the optimal parameters using a coarse-grained grid search Li et al. (2023a); Turner et al. (2023); Wang and Shu (2023).
@@ 3.3 Intervention to Enhance Distinct Trustworthiness Dimensions
Activation Intervention $\bm{v}_{ckpt\_{179}}$ 0.5799 0.5986 0.3524 0.2499 0.3914 0.2851 0.0600 0.5055 0.5390
$\bm{v}_{AmberChat}$ 0.5830 0.5958 0.3508 0.2519 0.3952 0.3352 0.0820 0.5055 0.5528
@@
Observation 2. Steering vectors derived from pre-training checkpoints and SFT model perform broadly comparable performance yet exhibit variations across various tasks. Table 1 shows that, compared to the steering vector extracted from AmberChat, the steering vector from the pre-training checkpoint (ckpt_179) guides the SFT model to exhibit more “truthfulness.” Moreover, it performs slightly better on ARC, ConfAIde, and StereoSet, while the opposite is true for other tasks. It is important to note that we only selected a single checkpoint from the pre-training process for experimentation, without undergoing fine-grained hyperparameter selection. Therefore, we believe these pre-training checkpoints hold significant untapped potential for aiding LLMs toward trustworthiness.
@@
Observation 3. Intervening in the model slightly impairs its general capabilities as a marginal cost for trustworthiness enhancement. We evaluate the model’s performance on four common benchmarks for general capabilities, where a trend of slight performance decline is observed after the intervention, as indicated in the “General Abilities” part of Tables 1 and 2. Additionally, we also observe the impact of the intervention strength $\alpha$ on the generative performance of the model. Taking ToxiGen as an example, Figure 3 illustrates the relationship between the proportion of toxic content generated by the model and perplexity as the intervention strength $\alpha$ increases. If we continuously increase the intervention strength, although the proportion of toxicity may continue to decline, the perplexity of the model correspondingly increases, manifesting as a tendency to produce meaningless repetitive content or gibberish.
Observation 4. When the quantity and quality of fine-tuning data are limited, activation intervention by steering vectors may be a more effective approach for the current task. We fine-tune the SFT model with positive QA pairs from the training set using both full-parameter fine-tuning and LoRA fine-tuning as a comparison, given that data in TruthfulQA naturally exists in the form of QA pairs. As shown in Table 1, the model fine-tuned with all parameters exhibits only minor improvements on TruthfulQA while experiencing a significant decline in general capabilities. Meanwhile, the fine-tuned model by LoRA demonstrates a noticeable decrease in TruthfulQA, though somewhat preserving performance in general capabilities.